Assuming that the equation f(Q, P1, P2, Y) = 10P1Q1 + 5Q1 - 2P2 - 4Y -18 =0


defines an implicit demand function Q1 = Q1(P1, P2, Y), find own price elasticity , cross elasticity and income elasticity at a point (P1, P2, Y) = (2, 1, 20)

1
Expert's answer
2022-01-23T15:55:57-0500

"f(Q_1P_1P_2,Y)= 10P_1Q_1+5Q_1-2P_2-4Y-18"


Where; (P1, P2, Y) = (2, 1, 20)


"Q_1= \frac{2P_2+4Y+18}{10P_1+5}" =4


"Own Price Elasticity=\frac{\Delta Q}{\Delta P}\times \frac{p}{Q}"

"=\frac{\Delta Q}{\Delta P_1}=10\times \frac{P}{Q}"


"10\times \frac{2}{4}="5

"Own Price Elasticity=\frac{\Delta Q}{\Delta P_2}\times \frac{p_2}{Q}"


Cross elasticity= "\frac{1}{4}\times -2= \frac{-2}{4}= -0.5"


"Income Elasticity =\frac{\Delta Q}{\Delta I}\times \frac{I}{Q}"

"=\frac{-4}{4}\times 20=-20"